OpenStudy (anonymous):

Solve the differential equation dy/dx = xy where y=10 when x=0. Give your answer in the form y=f(X).

myininaya (myininaya):

\[\frac{1}{y} dy= x dx\]

myininaya (myininaya):

integrate both sides

myininaya (myininaya):

\[\ln(y)=\frac{x^2}{2}+C\]

myininaya (myininaya):

\[y=e^{\frac{x^2}{2}+C}\]

myininaya (myininaya):

you can do the rest right?